Comparative figures for Apple and Google follow. $ millions Inventory Cost of sales Current Year $ 4,855 141,048 Apple One YeRequired 1 Required 2 Required 3 Compute inventory turnover for each company for the most recent two years shown. (Round yourRequired 1 Required 2 Required 3 Compute days sales in inventory for each company for the three years shown. (Round your ansRequired 1 Required 2 Required 3 In the current year, does (a) Apples and (b) Microsofts inventory turnover underperform or
Answer #1
1) Average Inventory = (beginning inventory + ending inventory)/2
Current Year One Year Ago
Apple = ( $2132+4855)/2 = ( $2349+2132)/2
= $ 3493.5 = $ 2240.5
Google = ( $268+749)/2 = ( $491+268)/2
= $ 508.5 = $ 379.5
Inventory Turnover Ratio = Cost of goods sold / average inventory
Current Year One Year Ago
Apple = $141048/3493.5 = $131376/2240.5
=40.4times =58.6 times
Google = $45583/508.5 = $35138/379.5
=89.6 times =92.6 times
2) Days sales in inventory = 365/ inventory turnover ratio
Current Year One Year Ago
Apple = 365 days /40.4 = 365 days /58.6
=9 days =6.2 days
Google = 365 days /89.6 = 365 days /92.6
=4.1 days =3.9 days
3) Apple outperformed
Google outperformed
Kindly give a positive rating if you are satisfied with this solution and please ask if you have any queries.

> where is two years prior? i think this is incorrect.

ariel friday Wed, Jan 19, 2022 5:30 PM